can you tell me which one of this two =to 0 and can you give me the whole work out how the answer came up to (4x+1) or 2x-7) thank you ,give me all of the work that came up with thwanswer so that I can understand how the answer got there so I know what i'm doing wrong. thank you
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

What is the original problem without the babbling?

Answer by stanbon(75887)

can you tell me which one of this two =to 0 and can you give me the whole work out how the answer came up to (4x+1) or (2x-7) thank you ,give me all of the work that came up with thwanswer so that I can understand how the answer got there so I know what i'm doing wrong.
----------------------
Neither 4x+1 or 2x-7 is zero, unless you say it is.
---
But, if 4x+1 = 0
then 4x = -1
and x = -1/4
-----------------
Similarly, if 2x-7 = 0
then 2x = 7
and x = 7/2
=====================
Cheers,
Stan H.